OpenStudy (anonymous):

Does anybody know a good website that goes through a step to step process about how to integrate ?

OpenStudy (anonymous):

I can help...and i'll give you examples to solve as well

OpenStudy (anonymous):

thank you but i meant like lectures, i have an exam coming up and i'm really confused about the whole unit

OpenStudy (anonymous):

it's part of first year university but my grade 12 teacher is making us do it now

OpenStudy (anonymous):

Is it just basic integration? So i'm guessing that's basic integration.

OpenStudy (anonymous):

it's basic, partial, area, and definite

OpenStudy (anonymous):

I'm not sure what you mean by partial because integration by parts is more than basic integration. You would have to know exponential functions and log functions as well as natural functions.

OpenStudy (anonymous):

natural log functions* and trig functions

OpenStudy (anonymous):

But anyway, I can explain integration step by step with examples then if it doesn't help, i'll give you a link

OpenStudy (anonymous):

okay thanks ! :)

OpenStudy (anonymous):

i have a link to my teacher's website if you want to check out what he's doing

OpenStudy (anonymous):

it's just not clear to me in words what he's doing and what method he applies

OpenStudy (anonymous):

ok sure, what's the link?

OpenStudy (anonymous):

http://sh.triton.net/math.htm the last calender for mhf4u the last few notes

OpenStudy (anonymous):

ok so I assume you have done integration of log/natural log functions, exponential functions and trig functions?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok so I opened the notes for May 29, what exactly don't you understand?

OpenStudy (anonymous):

okay i understand everything up till the dy/dx that really confuses me

OpenStudy (anonymous):

like i just dont understand what to do when i see that

OpenStudy (anonymous):

I don't think we have a dy/dx anywhere in integration by parts unless you are referring to du/dx? after substituting by u?

OpenStudy (anonymous):

yes typo :$ !

OpenStudy (anonymous):

so i just substitute it from there on or what?

OpenStudy (anonymous):

ok I'll start from basics. In Integration we can have two functions multiplied together, this makes things complicated. So we either use u substitution or integration by parts.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

U substitution is usually much simpler. Let me give you a couple of examples.

OpenStudy (anonymous):

okay that would be great

OpenStudy (anonymous):

Evaluate \[\int\limits_{}^{} (x^2 +1)/(x^3+3x) dx\]

OpenStudy (anonymous):

lett (x^3 + 3x) be u ?

OpenStudy (anonymous):

If you want to substitute first look at what you are trying to substitute and see if it will give you the other function after differentiation. In this case, u= x^3 + 3x because du=3(x^2+1)dx notice you're du almost looks like x^2 +1 which is the other function that we are trying to integrate?

OpenStudy (anonymous):

oh so the other function would be u ?

OpenStudy (anonymous):

\[u= x^3 +3x\] so once we have u, we differentiate it with respect to x. so that means \[du/dx= 3x^2 +1\] The reason why we do all this, is because we need to get rid of the dx, along with the other function, and just have one function that we can easily integrate.

OpenStudy (anonymous):

oh okay !

OpenStudy (anonymous):

I mean \[du/dx= 3x^2 + 3\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

So we multiply both sides of the equation to get du= 3(x^2 + 1) dx

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

do multipy

OpenStudy (anonymous):

But then since I want to get rid of \[(x^2 +1) dx\] I have to make du equal to \[(x^2 +1) dx\] This is because when we substituted x^3 + 3x with u, the integral looked like this \[\int\limits_{}^{} (x^3+1) dx/u\]

OpenStudy (anonymous):

why do we have to put the dx in ?

OpenStudy (anonymous):

Into the integral you mean?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Whenever you write an integral you will notice a dx dt dv, depending on what the function is. The dx,dt,dv or whatever it may be is part of the function that you are trying to integrate. So you can not ignore it.

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

So back to the example

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

I have to make my du similar to (x^3 +1) dx so that I can replace the whole thing with du and be left with this integral\[\int\limits_{}^{} du/u\]

OpenStudy (anonymous):

so is that the optimal ? like that's what format we want to reach ?

OpenStudy (anonymous):

yes that is what you want to have, and then you can integrate from here. The goal of substitution always, is to be left with a du and u. Could be \[\int\limits_{}^{} udu\] \[\int\limits_{}^{} du/u\] \[\int\limits_{}^{} u^2 du\] As long as we have a U and dU ONLY.

OpenStudy (anonymous):

okay ill write this down so i remember :P

OpenStudy (anonymous):

okay, so we had \[du= 3(x^2 +1)dx\] right?

OpenStudy (anonymous):

yupp

OpenStudy (anonymous):

So we divide both sides by 3 to get \[1/3 du= (x^2 + 1)dx\]

OpenStudy (anonymous):

okayy

OpenStudy (anonymous):

Now our du is exactly similar to the (x^2 + 1)dx in our integral, so we substitute this with 1/3 du to have \[\int\limits_{}^{} du/3u\] because it's 1/3 du divided by u

OpenStudy (anonymous):

ohh okay

OpenStudy (anonymous):

but then since 1/3 is a constant we can put it before the integral to have \[1/3 \int\limits_{}^{} du/u\]

OpenStudy (anonymous):

Constants ALWAYS factor out.

OpenStudy (anonymous):

oh okay do they equal zero

OpenStudy (anonymous):

no, you will multiply your constant to the value you get from your integral.

OpenStudy (anonymous):

ohh okay !

OpenStudy (anonymous):

Yea, so now we're on the last step, and this is when we integrate, remember this, \[\int\limits_{}^{} du/u = \ln(u) +C \] ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So the answer to our integral would be \[1/3 \int\limits_{}^{} du/u= 1/3 (\ln u)+ C \] and then we just plug in our initial value of u, which was x^2 + 3x to get \[\ln \left| (x^3 +3x \right| + C\] The absolute values just mean that ln x>0 so you may ignore them.

OpenStudy (anonymous):

oh okaay !

OpenStudy (anonymous):

so just as recap of what we did, we start by looking for the best function to substitute, the function whose derivative will be similar to the other function. Then we find du/dx. Then we equal du to the derivative of the function * dx. du=xdx for example. Then we replace the function xdx with du in the integral. and integrate our u with respect to du.

OpenStudy (anonymous):

How about you try one example of integration by u substitution?

OpenStudy (anonymous):

honestly thank you soooooo much !

OpenStudy (anonymous):

thanks so much but i have to get off now goodnight:)

OpenStudy (anonymous):

I mean I can give you the problem and solution and you can try it later?

OpenStudy (anonymous):

that would be great ! :)

OpenStudy (anonymous):

\[\int\limits_{}^{} 3x+6/(x^2 +4x -3) dx\]

OpenStudy (anonymous):

tyy !

OpenStudy (anonymous):

solution is \[(3/2) \ln \left| x^2 + 4x -3 \right| + C\]

OpenStudy (anonymous):

and your welcome. Good luck. Unfortunate we didn't get to do integration by parts. But good luck.

OpenStudy (anonymous):

thankyou :)